# Enabling Population-Level Parallelism in Tree-Based Genetic Programming for GPU Acceleration

### v5

@@ Abstract
Tree-based Genetic Programming (TGP) is a widely used evolutionary algorithm for tasks such as symbolic regression, classification, and robotic control. Due to the intensive computational demands of running TGP, GPU acceleration is crucial for achieving scalable performance. However, efficient GPU-based execution of TGP remains challenging, primarily due to three core issues: (1) the structural heterogeneity of program individuals, (2) the complexity of integrating multiple levels of parallelism, and (3) the incompatibility between high-performance CUDA execution and flexible Python-based environments. To address these issues, we propose EvoGP, a high-performance framework tailored for GPU acceleration of TGP via population-level parallel execution. First, EvoGP introduces a tensorized representation that encodes variable-sized trees into fixed-shape, memory-aligned arrays, enabling uniform memory access and parallel computation across diverse individuals. Second, EvoGP adopts an adaptive parallelism strategy that dynamically combines intra- and inter-individual parallelism based on dataset size, ensuring high GPU utilization across a broad spectrum of tasks. Third, EvoGP embeds custom CUDA kernels into the PyTorch runtime, achieving seamless integration with Python-based environments such as Gym, MuJoCo, Brax, and Genesis. Experiments show that EvoGP attains a peak throughput exceeding $10^{11}$ GPops/s, with speedups of up to $528\times$ over GPU-based TGP implementations and $18\times$ over the fastest CPU-based libraries, while maintaining comparable accuracy and improved scalability across large population sizes. EvoGP is open source and accessible at: https://github.com/EMI-Group/evogp.
@@ I Introduction
    We empirically evaluate the proposed framework, which attains a peak throughput exceeding $10^{11}$ GPops/s, with speedups of up to $528\times$ over GPU-based TGP implementations and $18\times$ over the fastest CPU-based libraries, while improving scalability for large population sizes.
@@ II-C Existing TGP Implementations
DEAP [27] and gplearn [28] are representative CPU-based libraries for genetic programming. DEAP is designed for extensibility and supports a wide range of evolutionary algorithms with customizable operators and multiprocessing, whereas gplearn focuses on symbolic regression and classification and employs joblib to parallelize evaluations across CPU cores. In contrast, Operon [29] is a modern CPU-based library that incorporates advanced parallel computing techniques to improve computational efficiency. Its C++ implementation exploits population-level parallelism through Intel Threading Building Blocks and data-level parallelism through SIMD vectorization, providing notable acceleration compared with earlier CPU-based GP libraries.
@@ III-B Tensorized Operations
where $s=\text{index}(k)$ is the starting position of the original subtree in $T_{\text{old}}$, and $e=s+\mathbf{n}_{\text{size}}^{\text{old}}(k)$ is the ending position of the original subtree in $T_{\text{old}}$. The tensor $\mathbf{n_{\text{size}}^{\text{updated}}}$ is computed by adding $\Delta n=\mathbf{n_{\text{size}}^{\text{new}}}[1]-\mathbf{n}_{\text{size}}^{\text{old}}[k]$ to the subtree sizes of all ancestor nodes of $k$:
@@ IV-A Overview of EvoGP
The overall system architecture of EvoGP is depicted in Fig.5, which illustrates its modular components and their interactions. Specifically, EvoGP is implemented in Python and leverages the PyTorch framework [44] for efficient tensor manipulation and GPU resource management. To optimize computational performance, EvoGP incorporates a suite of custom CUDA kernels for critical evolutionary operations, including tree initialization, mutation, crossover, inference, and fitness evaluation in symbolic regression (SR) tasks. To accommodate input datasets of varying sizes during SR evaluation, EvoGP adopts both data-parallel and hybrid-parallel execution strategies, as detailed in Section III-C. These CUDA kernels are integrated into the PyTorch runtime as custom operators, enabling seamless compatibility with the broader PyTorch ecosystem. This architecture combines the flexibility of Python-based development with the high-throughput capabilities of GPU acceleration, making EvoGP both user-friendly and computationally efficient.
@@ IV-C Multi-Output Support
Name
One Point
Cross.
One Point
Leaf Biased
Cross.
Subtree
Mut.
Hoist
Mut.
Single Point
Mut.
Multi-Point
Mut.
Insert
Mut.
Delete
Mut.
Single Const
Mut.
Multi-Const
Mut.
@@
SRGPU
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
@@ V-C Comparison with Existing TGP Implementations
To benchmark the performance of EvoGP, we selected several representative libraries for comparison. For CPU-based frameworks, three prominent TGP libraries were considered: DEAP [27], gplearn [28], and Operon [29]. For GPU-based counterparts, we included SRGPU [35]. In addition, PySR [58], a widely used library for symbolic regression, was incorporated as an additional reference. A supplementary comparison with another GPU library, TensorGP [34], is provided in Section S.I of the Supplementary Document. All benchmarks were performed on four symbolic regression tasks from the UCI Machine Learning Repository [59], as summarized in Table IV. Each experiment was executed for 100 generations. For clarity of visualization, we only present results of the runs that completed within one hour to ensure a consistent and practical comparative timeframe.
@@
Concrete Compressive Strength [62] 8 1030
Airfoil Self-Noise [63] 5 1503
@@
Population
Size
GPops/s by Dataset Size (Datapoints)
60 392 1030 1503
100 $6.37\times 10^{8}$ $2.97\times 10^{9}$ $9.54\times 10^{9}$ $1.18\times 10^{10}$
500 $1.59\times 10^{9}$ $8.76\times 10^{9}$ $2.35\times 10^{10}$ $3.18\times 10^{10}$
1,000 $4.21\times 10^{9}$ $2.23\times 10^{10}$ $4.20\times 10^{10}$ $5.14\times 10^{10}$
5,000 $1.03\times 10^{10}$ $5.74\times 10^{10}$ $6.59\times 10^{10}$ $8.26\times 10^{10}$
10,000 $1.29\times 10^{10}$ $6.79\times 10^{10}$ $7.36\times 10^{10}$ $8.26\times 10^{10}$
50,000 $3.23\times 10^{10}$ $9.78\times 10^{10}$ $6.68\times 10^{10}$ $8.14\times 10^{10}$
100,000 $3.68\times 10^{10}$ $1.00\times 10^{11}$ $8.15\times 10^{10}$ $9.15\times 10^{10}$
500,000 $3.33\times 10^{10}$ $9.22\times 10^{10}$ $7.96\times 10^{10}$ $8.91\times 10^{10}$
1,000,000 $3.98\times 10^{10}$ $9.90\times 10^{10}$ $7.80\times 10^{10}$ $8.67\times 10^{10}$
While wall-clock time provides a direct measure of speed, Langdon [64] introduced a metric that more precisely reflects the operational throughput of GP systems, termed GP operations per second (GPops/s). This metric represents the number of genetic program nodes evaluated per second and is defined as
$\text{GPops/s}=\frac{G\times P\times\bar{S}}{T},$
where $G$ denotes the number of generations, $P$ the population size, $\bar{S}$ the average program size measured by the number of nodes, and $T$ the total execution time in seconds.
@@
The GPops/s performance of EvoGP under different settings is presented in Table V. As evidenced, when the population size reaches 5,0005{,}000, the performance of EvoGP remains stable above $10^{10}$ GPops/s and peaks at $10^{11}$ GPops/s. Based on this throughput, we compared EvoGP with the fastest competing libraries. Fig. 9 and Fig. 10 present the GPops comparison between EvoGP and the fastest GPU-based library SRGPU and the fastest CPU-based library Operon. The heatmaps indicate EvoGP’s performance advantage. On smaller problems, EvoGP attains up to $528\times$ the GPops/s of SRGPU and $18\times$ that of Operon at a population size of $10^{6}$. On larger problems, EvoGP maintains this advantage and delivers $36\times$ the GPops/s of SRGPU and nearly $8\times$ that of Operon under the same population setting. A detailed breakdown of these results, including the exact GPops/s values for all compared libraries, is provided in Section S.II of the Supplementary Document.
@@ V-D Performance in Multiple Tasks with Scalable Populations
We evaluate the performance of EvoGP with different population sizes across three domains: symbolic regression, classification, and robotics control. For all tasks, we test population sizes of 200, 1,000, and 5,000. In symbolic regression, the experiments use the same benchmarks described in Section V-C and adopt MSE as the performance metric, where lower values indicate better accuracy. In classification, we use four datasets from the UCI repository [59]: Wine, Breast Cancer Wisconsin, Yeast, and Estimation of Obesity Levels. These datasets differ in the number of input features, output classes, and data instances, with detailed specifications summarized in Table VI. Performance is measured by classification accuracy, where higher values indicate better results. In robotics control, we use four environments from the Brax framework [48]: Swimmer, Hopper, Walker2d, and HalfCheetah. Table VII lists the observation and action dimensions of these environments. Performance is evaluated by the cumulative reward obtained during an episode, where higher rewards represent improved control effectiveness.
@@
   Concrete Compressive Strength 111.460 91.921 50.767
   Airfoil Self-Noise 28.081 20.240 16.739
@@
The results are shown in Fig. 11 and Table VIII. The figure presents task performance with respect to wall-clock time and illustrates the convergence behaviour for each population size, whereas summarizes the final performance of each population at the end of the corresponding runtime. A comparison of performance over a fixed number of generations is provided in Section S.III of the Supplementary Document.
@@
Across all three domains, the results indicate that larger populations yield lower MSE in regression, higher accuracy in classification, and higher cumulative rewards in control tasks. As shown in Table VIII, the improvement is most evident in the HalfCheetah control task, where increasing the population from 200 to 5,000 more than doubled the cumulative reward, and in the Daily Demand Forecasting regression task, where the final MSE decreased by over two orders of magnitude. These observations suggest that larger populations may contribute to improved GP performance by enabling more comprehensive explorations of the solution space.
@@ VI Conclusion
To further enhance evaluation efficiency, we proposed an adaptive parallelism strategy that dynamically fuses intra- and inter-individual parallelism based on dataset size. EvoGP seamlessly integrates with Python environments via custom CUDA operators within the PyTorch runtime, thereby allowing effortless deployment in popular platforms such as Gym and MuJoCo. Experimental results show that EvoGP attains a peak throughput exceeding $10^{11}$ GPops/s, with speedups of up to $528\times$ over GPU-based TGP implementations and $18\times$ over the fastest CPU-based libraries, while preserving solution quality and improving scalability for large populations.

### v6

@@ Abstract
Tree-based Genetic Programming (TGP) is a widely used evolutionary algorithm for tasks such as symbolic regression, classification, and robotic control. Due to the intensive computational demands of running TGP, GPU acceleration is crucial for achieving scalable performance. However, efficient GPU-based execution of TGP remains challenging, primarily due to three core issues: (1) the structural heterogeneity of program individuals, (2) the complexity of integrating multiple levels of parallelism, and (3) the incompatibility between high-performance CUDA execution and flexible Python-based environments. To address these issues, we propose EvoGP, a high-performance framework tailored for GPU acceleration of TGP via population-level parallel execution. First, EvoGP introduces a tensorized representation that encodes variable-sized trees into fixed-shape, memory-aligned arrays, enabling uniform memory access and parallel computation across diverse individuals. Second, EvoGP adopts an adaptive parallelism strategy that dynamically combines intra- and inter-individual parallelism based on dataset size, ensuring high GPU utilization across a broad spectrum of tasks. Third, EvoGP embeds custom CUDA kernels into the PyTorch runtime, achieving seamless integration with Python-based environments such as Gym, MuJoCo, Brax, and Genesis. Experimental results demonstrate that EvoGP achieves a peak throughput exceeding $10^{11}$ GPops/s. Specifcially, this performance represents a speedup of up to $304\times$ over existing GPU-based TGP implementations and $18\times$ over state-of-the-art CPU-based libraries. Furthermore, EvoGP maintains comparable accuracy and exhibits improved scalability across large population sizes. EvoGP is open source and accessible at https://github.com/EMI-Group/evogp.
@@ I Introduction
    We empirically evaluate the proposed EvoGP. Extensive experiments demonstrate that EvoGP attains a peak throughput exceeding $10^{11}$ GPops/s, achieves speedups of up to $304\times$ over GPU-based TGP implementations and $18\times$ over the fastest CPU-based libraries, and improves scalability for large population sizes.
@@ II-C Existing TGP Implementations
DEAP [19] and gplearn [59] are representative CPU-based libraries for genetic programming. DEAP is designed for extensibility and supports a wide range of evolutionary algorithms with customizable operators and multiprocessing, whereas gplearn focuses on symbolic regression and classification and employs joblib to parallelize evaluations across CPU cores. In contrast, Operon [9] is a modern CPU-based library that incorporates advanced parallel computing techniques to improve computational efficiency. Its C++ implementation exploits population-level parallelism through Taskflow [28] and data-level parallelism through SIMD vectorization, which provides notable acceleration compared with earlier CPU-based GP libraries.
@@ III-B Tensorized Operations
where $s=\text{index}(k)$ and $e=s+\mathbf{n}_{\text{size}}^{\text{old}}[k]$ denote the starting and ending positions, respectively, of the original subtree in $T_{\text{old}}$. The tensor $\mathbf{n_{\text{size}}^{\text{updated}}}$ is computed by adding $\Delta n=\mathbf{n_{\text{size}}^{\text{new}}}[1]-\mathbf{n}_{\text{size}}^{\text{old}}[k]$ to the subtree sizes of all ancestor nodes of $k$:
@@ IV-A Overview of EvoGP
The overall system architecture of EvoGP is depicted in Fig.5, which illustrates its modular components and their interactions. Specifically, EvoGP is implemented in Python and leverages the PyTorch framework [49] for efficient tensor manipulation and GPU resource management.
To optimize computational performance, EvoGP incorporates a suite of custom CUDA kernels for critical evolutionary operations, including tree initialization, mutation, crossover, inference, and fitness evaluation in symbolic regression (SR) tasks. To accommodate input datasets of varying sizes during SR evaluation, EvoGP adopts both data-parallel and hybrid-parallel execution strategies, as detailed in Section III-C. These CUDA kernels are integrated into the PyTorch runtime as custom operators, enabling seamless compatibility with the broader PyTorch ecosystem.
Complementing these kernel-level optimizations, EvoGP adopts a model for full-GPU residency, which ensures that data is not frequently transferred between the host and the device. By maintaining both the population and the evaluation context entirely on the GPU, EvoGP effectively eliminates the substantial communication overhead typical of traditional frameworks. This zero-copy design philosophy facilitates compatibility with modern reinforcement learning environments that are GPU-accelerated. Consequently, the architecture enables parallel simulations that are both efficient and end-to-end, thereby eliminating potential IO bottlenecks. Collectively, this design integrates the flexibility of Python-based development with the high-throughput capabilities of GPU acceleration, thereby ensuring that EvoGP is accessible and computationally efficient.
@@ IV-C Multi-Output Support
Name One Point Cross. One Point Leaf Biased Cross. Subtree Mut. Hoist Mut. Single Point Mut. Multi-Point Mut. Insert Mut. Delete Mut. Single Const Mut. Multi-Const Mut.
@@
SRGPU $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
@@ V-C Comparison with Existing TGP Implementations
To benchmark the performance of EvoGP, we selected several representative libraries for comparison. For CPU-based frameworks, we considered three established TGP libraries: DEAP [19], gplearn [59], and Operon [9]. Since Operon provides alternative mathematical backends that affect performance results, we utilized the Eve backend to enhance computational throughput. For GPU-based counterparts, we included SRGPU [69]. Furthermore, PySR [12] was incorporated as an additional reference. A supplementary comparison with the TensorGP GPU library [4] is provided in Section S.I of the Supplementary Document. All benchmarks were performed on four symbolic regression tasks collected from diverse sources. These tasks include Daily Demand Forecasting and Auto MPG from the UCI Machine Learning Repository [15], California Housing from scikit-learn [50], and Feynman I.9.18 from the Feynman Symbolic Regression Database [65]. Detailed specifications of these tasks are summarized in Table IV. Each experiment was executed for 100 generations. To ensure a consistent comparative timeframe, we only present results for the runs that completed within one hour.
@@
California Housing [29] 8 20,640
Feynman I.9.18 [65] 9 100,000
@@
Population Size GPops/s by Dataset Size (Datapoints)
60 392 20,640 100,000
100 $2.05\times 10^{8}$ $1.26\times 10^{9}$ $3.49\times 10^{10}$ $1.55\times 10^{11}$
500 $1.20\times 10^{9}$ $8.67\times 10^{9}$ $1.27\times 10^{11}$ $3.35\times 10^{11}$
1,000 $1.85\times 10^{9}$ $1.02\times 10^{10}$ $1.61\times 10^{11}$ $3.50\times 10^{11}$
5,000 $7.31\times 10^{9}$ $2.93\times 10^{10}$ $2.07\times 10^{11}$ $4.14\times 10^{11}$
10,000 $1.13\times 10^{10}$ $4.32\times 10^{10}$ $2.08\times 10^{11}$ $3.92\times 10^{11}$
50,000 $1.67\times 10^{10}$ $3.38\times 10^{10}$ $2.14\times 10^{11}$ $4.10\times 10^{11}$
100,000 $2.27\times 10^{10}$ $5.31\times 10^{10}$ $2.12\times 10^{11}$ $4.05\times 10^{11}$
500,000 $3.25\times 10^{10}$ $5.87\times 10^{10}$ $1.99\times 10^{11}$ $3.78\times 10^{11}$
1,000,000 $3.29\times 10^{10}$ $5.87\times 10^{10}$ $1.97\times 10^{11}$ $3.90\times 10^{11}$
While wall-clock time provides a direct measure of speed, Langdon [35] introduced a metric that more precisely reflects the operational throughput of GP systems, termed GP operations per second (GPops/s). The GPops/s metric is defined as the total number of genetic program primitive operations evaluated per second across all fitness cases, calculated as:
$\text{GPops/s}=\frac{G\times P\times\bar{S}\times D}{T},$
where $G$ denotes the number of generations, $P$ the population size, $\bar{S}$ the average program size measured by the number of nodes, $D$ the number of individual fitness cases (i.e., data points), and $T$ the total execution time in seconds.
@@
EvoGP achieves robust computational throughput, maintaining a magnitude of approximately $10^{10}$ GPops/s on small datasets and reaching $10^{11}$ GPops/s on medium-to-large datasets, as detailed in Table V. Based on this high throughput, we compared EvoGP against the leading GPU-based library, SRGPU, and the CPU-based library, Operon, as illustrated in Fig. 9 and Fig. 10. The results indicate that EvoGP holds a substantial performance advantage. Specifically, on smaller problems, EvoGP achieves peak speedups of $304\times$ relative to SRGPU and $18\times$ relative to Operon. In scenarios involving medium to large datasets, data-point parallelism saturates GPU resources, which naturally constrains the speedup magnitude. Nevertheless, EvoGP maintains a robust lead, thereby delivering at least a $3\times$ speedup over SRGPU. Against Operon, EvoGP sustains an approximately $10\times$ advantage, a margin that becomes particularly pronounced when the population size exceeds 1,0001{,}000. Although the hardware generational gap favors the GPU (14nm CPU vs. 8nm GPU), the order-of-magnitude speedups achieved by EvoGP remain substantial. Consequently, even with an estimated 40% gain from a more modern 7nm CPU architecture, the throughput advantage of EvoGP remains clearly evident. A detailed breakdown of these results, including the exact GPops/s values for all compared libraries, is provided in Section S.III of the Supplementary Document.
@@ V-D Performance in Multiple Tasks with Scalable Populations
We evaluate the performance of EvoGP across three domains: symbolic regression, classification, and robotics control. For all tasks, we test population sizes of 200, 1,000, and 5,000. In symbolic regression, the experiments utilize the benchmarks described in Section V-C and adopt MSE as the performance metric. Specifically, lower values indicate superior accuracy. In classification, we employ four datasets from the UCI repository [15]: Wine, Chronic Kidney Disease, Breast Cancer Wisconsin, and Yeast. These datasets exhibit heterogeneity in the number of input features, output classes, and data instances. Table VI summarizes these specifications. We measure performance via classification accuracy, where higher values denote better results. In robotics control, we utilize four environments from the Brax framework [20]: Swimmer, Hopper, Walker2d, and HalfCheetah. Table VII lists the observation and action dimensions of these environments. Performance is evaluated by the cumulative reward obtained during an episode. Consequently, higher rewards represent improved control effectiveness.
@@
   California Housing 0.533 0.451 0.426
   Feynman I.9.18 2.207e-3 9.274e-4 7.239e-4
@@
The results are shown in Fig. 11 and Table VIII. The figure presents task performance with respect to wall-clock time and illustrates the convergence behaviour for each population size, whereas summarizes the final performance of each population at the end of the corresponding runtime. A comparison of performance over a fixed number of generations is provided in Section S.IV of the Supplementary Document.
@@
Across all three domains, the results indicate that larger populations yield lower MSE in regression, higher accuracy in classification, and higher cumulative rewards in control tasks. As shown in Table VIII, the improvement is most evident in the HalfCheetah control task, where increasing the population from 200 to 5,000 more than doubled the cumulative reward, and in the Daily Demand Forecasting regression task, where the final MSE decreased by over two orders of magnitude. Similarly, in the Yeast classification task, increasing the population size enabled EvoGP to surpass the standard accuracy of Neural Networks ($56.87\%$) and closely approach the best accuracy reported in the UCI repository benchmarks ($62.80\%$, achieved by Random Forest). These observations suggest that larger populations may contribute to improved GP performance by enabling more comprehensive explorations of the solution space.
@@ VI Conclusion
To further enhance evaluation efficiency, we proposed an adaptive parallelism strategy that dynamically fuses intra- and inter-individual parallelism based on dataset size. EvoGP seamlessly integrates with Python environments via custom CUDA operators within the PyTorch runtime, thereby allowing effortless deployment in popular platforms such as Gym and MuJoCo. Experimental results showdemonstrate that EvoGP attains a peak throughput exceeding $10^{11}$ GPops/s, withwhich corresponds to speedups of up to $304\times$ over GPU-based TGP implementations and $18\times$ over the fastest CPU-based libraries. Furthermore, whileEvoGP preservingpreserves solution quality and improvingimproves scalability for large populations.